I don't need 4 closing entries. I need a list of current assets, a list of current liabilities, a working capital, and a current ratio.The following is the adjusted trial balance for Harrahs Casino as of December 31, 2016. Harrahs Casino Trial Balance Decemb
Answer #1

Current Assets -

1. Cash 3,000

2. Account Receivable 600

3. Prepaid Rent 900

Total Current Asset = 4,500

---

Current Liabilities -

1. Accounts Payable 1,500

2. Salaries Payable 3,000

Total Current Liabilities = 4,500

---

Working Capital = Current Asset - Current Liabilities = 4,500 - 4,500 = 0.

---

Current Ratio = \frac{Current Asset}{Current Liabilities}= \frac{4,500}{4,500}= 1

Note : Poker Chips ,Table and equipment are Non-Current assets.